# Post-detection inference for sequential changepoint localization

### v1

@@ 3.3 A concrete example: Gaussian mean shift problem
We consider the most common changepoint model, where the pre-change distribution is $F_{0}$ is $N(\theta_{0},1)$ and the post-change model is $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$.
In this setting, we have the following example of a $1-\beta$ confidence sequence (Howard et al.,, 2021) for $\theta_{1}$
@@
We fix $t\in\{1,\cdots,\tau\}$. Note that we can consider $\theta^{\prime\prime}=\theta_{0}$ and $G(x,\theta)=x+\theta-\theta_{0}$. Therefore, only need to generate the sequence $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(\theta_{0},1)$ for $j=1,\cdots,B$. For $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$,
$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\epsilon_{n}^{j}+\theta_{0},\text{ if }n<t\\ \epsilon_{n}^{j}+\theta,\text{ otherwise, }\end{cases}$ (3.14)
where $\epsilon_{n}^{j}:=X_{n}^{j}-\theta_{0}\sim N(0,1)$. We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (2.8), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Let, $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$, where $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$. Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark 3.3) and they are finite. So, we can have $L=\infty$.
If $t_{2}^{j}<t$, we have $V_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t,V_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
$V_{t,\infty}^{j}$ $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}(\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))}{L_{t}(\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j}+\theta;Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))}$
@@
$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta_{t^{\prime}}^{i});Y_{1}^{t}(X_{1}^{j},\theta_{t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta_{t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$ (3.15)
@@
$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta-\theta_{0})\right)^{2}+\text{constant (wrt }\theta),\text{ if }i<t,\\ \text{constant (wrt }\theta),\text{ if }i=t,\\ \displaystyle\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}+\theta-\theta_{0}\right)^{2}+\text{constant (wrt }\theta),\text{ if }i>t,\end{cases}$
over $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. Note that $\theta_{t^{\prime}}^{i}$ has a simple closed-form expression since the above equation is a quadratic function of $\theta$. And hence, $V_{t}^{j}$ (3.3) can be computed easily.
@@
For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$. Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{\frac{1}{2}(\theta-\theta_{0})\times\sum_{i=j}^{\tau}(2X_{i}-\theta_{0}-\theta)\}$ increses as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Similarly, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Also, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\neq\theta_{0}\}$, if we employ Gaussian weights (it is conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution. One can easily check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section 5.2.
@@
    The coupling function $G$ is available (which is always the case when distributions are continuous).
@@
    If we employ the likelihood-rato-based test statistic (2.8), then we have to be able to maximize the function $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).
@@ 4 Composite pre- and post-change
We now generate $N$ many i.i.d. sequences under $F_{\theta_{0}^{*}}$ until a change is detected using $\mathcal{A}$, recording the detection times as $\tau_{1},\cdots,\tau_{N}$, for some $N\in\mathbb{N}$, that will help us estimate $\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$, which is a lower bound on $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$. If $\mathcal{A}$ controls PFA at a known level $\delta\in(0,1)$, we can avoid this step and obtain a conditional/unconditional coverage guarantee without any such assumption.
Since both pre- and post-change parameters are unknown, for each $t$, we estimate them based on $X_{1},\cdots,X_{t-1}$ and $X_{t},\cdots,X_{\tau}$ respectively. However, as noted earlier, the sample size used for the estimates might be too small to yield reliable estimates. Therefore, instead of point estimates, we rely on interval estimates in order to achieve a predetermined control over the error.
Suppose that we have some confidence interval procedure $\operatorname{CI}(X_{1},\cdots,X_{n};1-c)$, which constructs $1-c$ confidence interval using $X_{1},\cdots,X_{n}$ for $\theta_{0}$ and a confidence sequence procedure $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n}$, which constructs $1-c$ confidence sequence using $\{X_{n}\}_{n}$ for $\theta_{1}$. If $\mathcal{A}$ controls PFA, we can simply construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta)$ for $\theta_{1}$ having $1-\beta$ coverage and $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma)$ for $\theta_{0}$ having $1-\gamma$ coverage. Otherwise, we construct $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})$ and $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, where $r_{t}:=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$. Despite $r_{t}$ being random, it is based on independent simulations, and so it turns out that marginalizing over $r_{t}$ yields that the CS and CI have coverage probabilities $1-\beta\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$ and $1-\gamma\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$ respectively; we formalize this more explicitly in the proof of Theorem 4.2.
@@
Our confidence set is the collection of all $t$ before $\tau$, for which we fail to reject $\mathcal{H}_{0,t}$ at level $\alpha\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$,
@@
where $\hat{\alpha}_{t}:=\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$. If $\mathcal{A}$ controls PFA at some known level $\delta\in(0,1)$ , we have the following confidence set, which is simpler (i.e., does not require $\tau_{i}$’s):
@@
Algorithm 4 contains an overview of the method.
Note that, unless $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ are finite sets, it can be computationally challenging to approximate the threshold $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$. However, in many cases, we can derive a suitable upper bound of this threshold, which is computationally and statistically efficient. We provide the details in Section 4.2. We offer a concrete example of the Gaussian mean shift problem in Section 4.3.

### v2

@@ 3.3 A concrete example: Gaussian mean shift problem
Let the pre-change distribution $F_{0}$ be $N(\theta_{0},1)$ and the post-change model be $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$.
In this setting, we have the following example of a $1-\beta$ confidence sequence [9] for $\theta_{1}$
@@
We fix $t\in\{1,\cdots,\tau\}$. Note that we can consider $\theta^{\prime\prime}=\theta_{0}$ and $G(x,\theta)=x+\theta-\theta_{0}$. Therefore, only need to generate $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(\theta_{0},1)$ for $j=1,\cdots,B$. For $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$,
$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\epsilon_{n}^{j}+\theta_{0},\text{ if }n<t,\\ \epsilon_{n}^{j}+\theta,\text{ otherwise, }\end{cases}$ (3.14)
where $\epsilon_{n}^{j}:=X_{n}^{j}-\theta_{0}\sim N(0,1)$. We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (2.8), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Let, $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$. Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark 3.3) and they are finite. So, we can have $L=\infty$.
If $t_{2}^{j}<t$, we have $V_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t$, it can be simplified in our setting as below:
$V_{t,\infty}^{j}$ $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}\left(\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-i+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}{L_{t}\left(\frac{\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j}}{t^{\prime}-t+1}+\theta;Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}$
@@
$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta_{t^{\prime}}^{i});Y_{1}^{t}(X_{1}^{j},\theta_{t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta_{t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$
@@
$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{(t-i)(\theta-\theta_{0})}{t^{\prime}-i+1}\right)^{2}+\text{constant},\text{ if }i<t,\\ \text{constant},\text{ if }i=t,\\ \displaystyle\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}+\theta-\theta_{0}\right)^{2}+\text{ constant},\text{ if }i>t,\end{cases}$
over $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. Here, constant refers to the terms that do not involve $\theta$. Note that $\theta_{t^{\prime}}^{i}$ has a simple closed-form expression since the above equation is a quadratic function of $\theta$. And hence, $V_{t,\infty}^{j}$ can be computed easily.
@@
For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$. Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{\frac{1}{2}(\theta-\theta_{0})\times\sum_{i=j}^{\tau}(2X_{i}-\theta_{0}-\theta)\}$ increases as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Similarly, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Also, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\neq\theta_{0}\}$, if we employ Gaussian weights (it is conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution. One can check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section 5.2.
@@
    The coupling function $G$ is available (which is always the case when distributions are continuous and in the case of discrete distributions, even if a suitable $G$ is not directly available, our method may still be implementable, as demonstrated in Section B.1 for the Poisson distribution).
@@
    If we employ the likelihood-ratio-based test statistic (2.8), then we must be able to maximize $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).
@@ 4 Composite pre- and post-change
We now generate $N$ many i.i.d. sequences under $F_{\theta_{0}^{*}}$ until a change is detected using $\mathcal{A}$, recording the detection times as $\tau_{1},\cdots,\tau_{N}$, for some $N\in\mathbb{N}$, that will help us estimate $\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$, which is a lower bound on $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$. If $\mathcal{A}$ controls PFA at a known level $\delta\in(0,1)$, we can avoid this step and still obtain conditional or unconditional coverage.
Since both pre- and post-change parameters are unknown, for each $t$, we estimate them based on $X_{1},\cdots,X_{t-1}$ and $X_{t},\cdots,X_{\tau}$ respectively. Instead of point estimates, we rely on interval estimates in order to achieve a predetermined control over the error.
Suppose that we have some confidence interval procedure $\operatorname{CI}(X_{1},\cdots,X_{n};1-c)$, which constructs $1-c$ confidence interval using $X_{1},\cdots,X_{n}$ for $\theta_{0}$ and a confidence sequence procedure $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n}$, which constructs $1-c$ confidence sequence using $\{X_{n}\}_{n}$ for $\theta_{1}$. If $\mathcal{A}$ controls PFA, we can simply construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta)$ for $\theta_{1}$ having $1-\beta$ coverage and $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma)$ for $\theta_{0}$ having $1-\gamma$ coverage. Otherwise, we construct $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})$ and $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, where $r_{t}:=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$. Despite $r_{t}$ being random, it is based on independent simulations, and so it turns out that marginalizing over $r_{t}$ yields that the CS and CI have coverage probabilities $1-\beta\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$ and $1-\gamma\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$ respectively; we formalize this more explicitly in the proof of Theorem 4.2.
@@
Our confidence set is the set of $t\leq\tau$ for which we fail to reject $\mathcal{H}_{0,t}$ at level $\alpha\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$,
@@
where $\hat{\alpha}_{t}:=\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$. If $\mathcal{A}$ controls PFA at some known level $\delta\in(0,1)$ , we have the following confidence set, which is simpler (i.e., does not require $\tau_{i}$’s):
@@
Algorithm 4 contains an overview. As before, even if $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ are uncountable, one can approximate the threshold $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ by discretizing the parameter space into a grid and evaluating the quantile for values of $(\theta,\theta^{\prime})$ on this grid, taking their maximum. Additionally, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient, and exact (i.e., does not require any approximation). We provide the details in Section 4.2. Additionally, we illustrate concrete examples of the Gaussian mean shift problem in Section 4.3 and the Poisson rate change problem in Appendix B.